OpenStudy (anonymous):

4-3x/2 > x-2/3 I got 32/19 and 7/4... but the answer key says x <16/11

OpenStudy (anonymous):

is this \(4-\frac{3}{2}x>x-\frac{3}{2}\) ?

OpenStudy (anonymous):

2 covers the whole equation on the left side and 3 on the right side

OpenStudy (anonymous):

ok let me try again

OpenStudy (anonymous):

\[\frac{4-3x}{2}>\frac{x-2}{3}\]

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

multiply both sides by 6 get \[3(4-3x)>2(x-2)\] then \[12-9x>2x-4\]

OpenStudy (anonymous):

add \(9x\) get \[12>11-4\]add 4 get \[16>11x\]divide by 11 gives your answer

OpenStudy (anonymous):

sorry i meant add \(9x\) get \(12>11x-4\)

OpenStudy (anonymous):

in multiplying by 6. How did you get the 3 on the left side and 2 on the right side?